# Localisation of Regularised and Multiview Support Vector Machine Learning

### v3

@@ Abstract.
We prove a few representer theorems for a localised version of the regularised and multiview support vector machine learning problem introduced by H.Q. Minh, L. Bazzani, and V. Murino, Journal of Machine Learning Research, 17(2016) 1–72, that involves operator valued positive semidefinite kernels and their reproducing kernel Hilbert spaces. The results concern general cases when convex or nonconvex loss functions and finite or infinite dimensional input spaces are considered. We show that the general framework allows infinite dimensional input spaces and nonconvex loss functions for some special cases, in particular in case the loss functions are Gâteaux differentiable. Detailed calculations are provided for the exponential least squares loss functions that leads to partially nonlinear equations for which a particular unconstrained potential reduction Newton’s approximation method can be used.
@@ 1. Introduction
Representer theorems are of a special interest in machine learning due to the fact that they reduce the problem of finding a minimiser for the learning map to the vector space spanned by the kernel functions, or operators, at the labeled and unlabeled input data. For classical versions of representer theorem, we recommend the monographs of [56] and [60]. There is a large literature on generalised representer theorems but in this article we refer to the unifying framework in vector valued reproducing kernel Hilbert spaces for semisupervised, manifold regularised and multiview machine learning, as investigated by [48] and the vast literature cited there.
@@
There are a few reasons that motivate the localised versions that we consider in this article. Firstly, for some of the labeled input points, in the multivariable case, some of the components of the labels (properties) may be missing or some additional components of the labels may be necessary in order to allow reliable information. This means that the underlying vector spaces of labels may have different dimensions and hence, making the vector spaces of the labels depend on the input points solves this obstruction. Secondly, when the input set $X$ shows a certain homogeneity, the localised version may not be needed but, when this set is more heterogeneous, the localised version brings the necessary flexibility. To be more precise, let us imagine the following scenario. The input set $X$ is a finite union of sets $X_{i}$, $i=1,\ldots,N$, where each $X_{i}$ shows homogeneity. By homogeneity we mean that, for each $i$, the properties (labels) associated to $x\in X_{i}$ are of the same type, in particular they live in the same Hilbert space ${\mathcal{W}}_{i}$. However, for different $i\neq j$, where $i,j\in\{1,\ldots,N\}$, the properties (labels) of the points $x\in X_{i}$ when compared to the properties of the points $x\in X_{j}$ may be different, meaning that the Hilbert spaces of labels ${\mathcal{Y}}_{i}$ and ${\mathcal{Y}}_{j}$ should be different. For example, if the set $X$ is the collection of all the cells of a body, we can see $X$ as the union of all its organs $X_{i}$, for $i=1,\ldots,N$, and it is clear that due to special functions that different organs have in the body, the properties (labels) of cells in different organs are generally of different type. Moreover, the learning function, which is calculated in terms of loss functions, may require different loss functions for different points that have different types of labels. For example, in the scenario of the body viewed as the union of its organs, comparison of properties of different cells should be performed differently for cells belonging to different organs. This justifies the dependence of the loss functions of the input points as well. As a further research project, the localised version of the machine learning problem might be used for investigating the process of spreading a malignant tumor (metastasis) in different organs of a body, provided that real data will be available and appropriate mathematical models will be obtained.
@@
The direction of semisupervised learning we follow here is manifold regularisation, cf. [7], [13], and [47], which attempts to learn the geometry of the input space by exploiting the given unlabeled data. The latter two papers are recent generalisations of the original scalar version of manifold regularisation of [7] to the vector valued setting. In [13], a vector valued version of the graph Laplacian $L$ is used, and in [47], $L$ is a general symmetric, positive operator, including the graph Laplacian. The vector valued setting allows one to capture possible dependencies between output variables by the use of, for example, an output graph Laplacian. For a comprehensive discussion on semisupervised learning and a thorough comparison with supervised and unsupervised learning, see the collection [17].
@@
Because reproducing kernel Hilbert spaces make an essential ingredient in these representer type theorems, some historical considerations are in order. The classical article of [2] provides an abstract formalisation of scalar reproducing kernel Hilbert spaces of many previous investigations and applications related to spaces of analytic functions, partial differential equations, and harmonic analysis, as of [45], [8], [49], [11], [25], [36], and [37]. From the point of view of probability theory, [35] investigated linearisations, or feature spaces, associated to scalar positive semidefinite kernels. An equivalent formulation of reproducing kernel Hilbert spaces by Hilbert spaces continuously embeded in a quasicomplete locally convex space was investigated by [58]. Then, scalar valued reproducing kernel Hilbert spaces found many applications in machine learning, see [60] and [56] for a comprehensive list of literature in this direction.
@@
Motivated by problems in operator theory and operator algebras, operator valued positive semidefinite kernels and either their linearisations (Kolmogorov decompositions) or their reproducing kernel Hilbert spaces have been considered by [62], [52], [51], [22], and [18], to cite a few. Investigations of operator valued positive Hermitian kernels that yield Krein spaces have been performed by [19], [20]. More general operator valued positive semidefinite kernels that yield reproducing kernel VH-spaces have been considered by [24] and [3], [4], and [5]. To make things more precise, recall that a vector space ${\mathcal{S}}$ endowed with a map ${\mathcal{S}}\ni s\mapsto s^{*}\in{\mathcal{S}}$ which is conjugate linear and involutive, in the sense that $(s^{*})^{*}=s$ for all $s\in{\mathcal{S}}$, is called a *-vector space. If, in addition, a convex cone ${\mathcal{S}}^{+}$ on ${\mathcal{S}}$ is specified such that for any $s\in{\mathcal{S}}^{+}$ we have $s=s^{*}$, we call ${\mathcal{S}}$ an ordered *-space and, when a complete locally convex topology on ${\mathcal{S}}$ is specified and which is related in a certain fashion with the convex cone ${\mathcal{S}}^{+}$, one calls ${\mathcal{S}}$ an admissible space. The concept of admissible space is a generalisation of the concept of $C^{*}$-algebra which is a mainstream domain in functional analysis, partly due to its strong connections with quantum theories, e.g. see [21] and [10]. Admissible spaces ${\mathcal{S}}$ are then used to consider gramians $[\cdot,\cdot]\colon\mathcal{X}\times{\mathcal{S}}\rightarrow{\mathcal{S}}$, for some vector space $\mathcal{X}$, which are vector valued inner products and that induce a certain topology on $\mathcal{X}$. If this topology is complete then we call $\mathcal{X}$ a VH-space (Vector Hilbert space). The concept of VH-space is a generalisation of the concept of Hilbert modules over $C^{*}$-algebras, e.g. see [40], but it appeared independently and related to problems in probability theory, see [41], [42], and [43]. In case we have a positive semidefinite kernel $K\colon X\times X\rightarrow{\mathcal{S}}$, for some admissible space ${\mathcal{S}}$, one can define a reproducing kernel VH-space as a generalisation of the classical reproducing kernel Hilbert space. For details and many examples see, for example, [4].
@@
We briefly describe the contents of this article. In order to introduce the localised version of the semisupervised, regularised, manifold regularised, and multiview learning problem, in Subsection 2.1 we firstly consider operator valued positive semidefinite kernels for which the entries are localised by a bundle of Hilbert spaces. For these kernels, we show how their reproducing kernel Hilbert spaces are constructed, their relation to the linearisations (Kolmogorov decompositions, feature spaces) and basic properties. Although we have been inspired by the approach in [19], we provide in the appendices proofs for all the statements we make, since in this form, they cannot be found elsewhere. In [52], reproducing kernel Hilbert spaces are considered for kernels over bundles of locally convex spaces and this might be a very interesting research problem in case applications to machine learning theory might be found.
@@
Further on, in Subsection 3.2 we consider real underlying Hilbert spaces and loss functions that are Gâteaux differentiable and then, in Theorem 3.7 we show that, under this additional assumption, a general representer theorem can be obtained for infinite dimensional input spaces. In Subsection 3.3 we work out the details for the loss functions defined by the least squares which leads to a linear problem in terms of the unkown coefficients, as in [48] while in Subsection 3.4 we work out the details for the loss functions defined by the exponential least squares which lead to a mixed problem, linear and nonlinear. Finally, in Subsection 3.5 we tackle algorithms to obtain approximations of the solutions for the latter machine learning problem by damped Newton approximation methods. Since in the framework of systems of nonlinear equations only local solutions are generally available, much research remains to be done.
@@ 2. A General Semisupervised Regularised and Multiview Machine Learning Problem
Let $X$ be a nonempty set and $\boldsymbol{H}=\{{\mathcal{H}}_{x}\}_{x\in X}$ a bundle of Hilbert spaces over the field ${\mathbb{F}}$, that is either ${\mathbb{R}}$ or ${\mathbb{C}}$, with base $X$, that is, ${\mathcal{H}}_{x}$ is a Hilbert space over ${\mathbb{F}}$, for any $x\in X$. In order to avoid confusion, let us note that the Hilbert spaces ${\mathcal{H}}_{x}$ are actually tagged Hilbert spaces, meaning for $x\neq y$ the spaces ${\mathcal{H}}_{x}$ and ${\mathcal{H}}_{y}$ are disjoint. An $\boldsymbol{H}$-operator valued kernel $K$ is a mapping defined on $X\times X$ such that $K(y,x)\in{\mathcal{B}}({\mathcal{H}}_{x},{\mathcal{H}}_{y})$ for all $x,y\in X$. Here and throughout this article, if ${\mathcal{H}}$ and ${\mathcal{K}}$ are two Hilbert spaces over the same filed ${\mathbb{F}}$, we denote by ${\mathcal{B}}({\mathcal{H}},{\mathcal{K}})$ the Banach space of all linear and bounded operators $T\colon{\mathcal{H}}\rightarrow{\mathcal{K}}$, endowed with the operator norm. We denote by $\mathfrak{K}(X;\boldsymbol{H})$ the collection of all $\boldsymbol{H}$-operator valued kernels on $X$ and it is clear that $\mathfrak{K}(X;\boldsymbol{H})$ is a vector space over ${\mathbb{F}}$.
@@
In particular, since $\boldsymbol{H}$ consists of tagged Hilbert spaces this means that if either $h\in{\mathcal{H}}_{x}$ or $k\in{\mathcal{H}}_{y}$ are not null and $x\neq y$ then $\widehat{h}\neq\widehat{k}$. However, if $h=0_{{\mathcal{H}}_{x}}$ and $k=0_{{\mathcal{H}}_{y}}$ with $x\neq y$, then $\widehat{h}=\widehat{k}=0_{{\mathcal{F}}(X;\boldsymbol{H})}$. Also, for any $f\in{\mathcal{F}}(X;\boldsymbol{H})$ we have
@@
The following theorem is a general version of some classical results, e.g. [35], [51], [22]. This is also a special case of [19]. We include its proof in Appendix A.
@@
Also, it is worth mentioning that by (rk2), for each $x\in X$, we actually have a bounded linear operator $K_{x}\colon{\mathcal{H}}_{x}\rightarrow{\mathcal{R}}$ defined by $K_{x}h:=K(\cdot,x)h$, for all $h\in{\mathcal{H}}_{x}$. This operator is bounded, as proven in (4.1). The following result is a generalisation of Moore-Aronszajn Theorem, [49], [2], [46], [16], [15], [47], [34]. Also, it is a special case of [19] and it is proven in Appendix B. In Appendix C we present a more direct construction of the reproducing kernel Hilbert space induced by an operator valued positive semidefinite kernel.
@@
One of the most important property of a reproducing kernel Hilbert space consists in the fact that, as a function space, its topology makes continuous all evaluation operators, see the proof in Appendix D.
@@
Secondly, the reproducing kernel Hilbert space ${\mathcal{H}}_{K}$ is associated to a vector valued positive semidefinite kernel for several reasons, but mainly because this is related to the multiview learning, cf. [23], [46], [59], [55], [47], [44], [34], [48]. Here we consider localised versions of these operator valued reproducing kernel Hilbert spaces.
@@
$V(y,z)=\frac{1}{1+\exp(y-z)},\quad y,z\in{\mathbb{R}}.$
@@
The third term in (2.15) combines vactor valued manifold regularisation, cf. [47], with multiview regularisation, cf. [55] and [61]. The parameter $\gamma_{I}$ may be taken $1$, without loss of generality, since it can be absorbed in $M$. Following [48], the operator multiview regularisation term $\gamma_{I}\langle\boldsymbol{f},M\boldsymbol{f}\rangle_{\boldsymbol{W}^{l+u}}$ is decomposed as
@@
This example follows essentially [48] for a within-view manifold regularisation via multiview graph Laplacians in support vector machine learning, cf. [61]. For manifold regularisation, a data adjaceny graph is defined in such a way that the entries measure the similarity or closeness of pairs of inputs. Given an undirected graph $G=({\mathcal{V}},E)$, where the vertices are ${\mathcal{V}}=\{1,\ldots,n\}$ and edges are simply pairs $(j,k)$, assume that for each edge $(j,k)\in E$ there is a weight $w_{j,k}$, and to each edge $(j,k)\not\in E$ we let $w_{j,k}=0$, in such a way that the weight matrix $W=[w_{j,k}]$ is Hermitian and nonnegative (positive semidefinite). Letting $v_{j,j}=\sum_{k=1}^{n}w_{j,k}$ and $v_{j,k}=0$ if $j\neq k$, we make the diagonal matrix $V=[v_{j,k}]$. We work under the assumptions that $v_{j,j}>0$ for all $j=1,\ldots,n$. Then the graph Laplacian matrix is $L:=V-W$, which is positive semidefinite. Sometimes it is useful to work with the normalised graph Laplacian $\widetilde{L}:=V^{-1/2}LV^{-1/2}$. For example, when each vertex $j$ is associated to a vector $h_{j}\in{\mathbb{R}}^{d}$, we can use the Gaussian weights
@@
In order to simplify the complexity of calculations, cf. [61], for most of the edges $(j,k)$ we take $w_{j,k}=0$ and only for neighbouring $(j,k)$, that is, $\|h_{j}-h_{k}\|\leq\epsilon$, for some integer $\epsilon>0$, we define the weights by (2.17).
@@
$\langle\mathbf{f},M_{W}\mathbf{f}\rangle_{{\mathcal{W}}^{u+l}}=\sum_{i=1}^{m}\sum_{j,k=1,\ j<k}w_{j,k}^{i}\|f^{i}(x_{j})-f^{i}(x_{k})\|^{2}_{\mathcal{Y}},$
@@
$E_{\boldsymbol{C},\boldsymbol{x}}f=\left(C_{x_{1}}K_{x_{1}}^{*}f,\dots,C_{x_{l}}K_{x_{l}}^{*}f\right),\quad f\in{\mathcal{H}}_{K}.$ (2.21)

### v4

@@ Abstract.
We prove some representer theorems for a localised version of a semisupervised, manifold regularised and multiview support vector machine learning problem introduced by H.Q. Minh, L. Bazzani, and V. Murino, Journal of Machine Learning Research, 17(2016) 1–72, that involves operator valued positive semidefinite kernels and their reproducing kernel Hilbert spaces. The results concern general cases when convex or nonconvex loss functions and finite or infinite dimensional underlying Hilbert spaces are considered. We show that the general framework allows infinite dimensional Hilbert spaces and nonconvex loss functions for some special cases, in particular in case the loss functions are Gâteaux differentiable. Detailed calculations are provided for the exponential least squares loss functions that lead to systems of partially nonlinear equations for which a particular different types of Newton’s approximation methods based on the interior point method can be used. Some numerical experiments are performed on a toy model that illustrate the tractability of the methods that we propose.
@@ 1. Introduction
Representer theorems are of a special interest in machine learning due to the fact that they reduce the problem of finding a minimiser for the learning map to the vector space spanned by the kernel functions, or operators, at the labeled and unlabeled input data. For classical versions of representer theorem, we recommend the monographs [61] and [66]. There is a large literature on generalised representer theorems but in this article we refer to the unifying framework in vector valued reproducing kernel Hilbert spaces for semisupervised, manifold regularised and multiview machine learning, as investigated by [51] and the vast literature cited there.
@@
There are a few reasons that motivate the localised versions that we consider in this article. Firstly, for some of the labeled input points, in the multivariable case, some of the components of the labels (properties) may be missing or some additional components of the labels may be necessary in order to allow reliable information. This means that the underlying vector spaces of labels may have different dimensions and hence, making the vector spaces of the labels depend on the input points solves this obstruction. Secondly, when the input set $X$ shows a certain homogeneity, the localised version may not be needed but, when this set is more heterogeneous, the localised version brings the necessary flexibility. To be more precise, let us imagine the following scenario. The input set $X$ is a finite union of sets $X_{i}$, $i=1,\ldots,N$, where each $X_{i}$ shows homogeneity. By homogeneity we mean that, for each $i$, the properties (labels) associated to $x\in X_{i}$ are of the same type, in particular they live in the same Hilbert space ${\mathcal{Y}}_{i}$. However, for different $i\neq j$, where $i,j\in\{1,\ldots,N\}$, the properties (labels) of the points $x\in X_{i}$ when compared to the properties of the points $x\in X_{j}$ may be different, meaning that the Hilbert spaces of labels ${\mathcal{Y}}_{i}$ and ${\mathcal{Y}}_{j}$ should be different. For example, if the set $X$ is the collection of all the cells of a body, we can see $X$ as the union of all its organs $X_{i}$, for $i=1,\ldots,N$, and it is clear that due to special functions that different organs have in the body, the properties (labels) of cells in different organs are generally of different type. Moreover, the learning function, which is calculated in terms of loss functions, may require different loss functions for different points that have different types of labels. For example, in the scenario of the body viewed as the union of its organs, comparison of properties of different cells should be performed differently for cells belonging to different organs. This justifies the dependence of the loss functions of the input points as well. As a further research project, the localised version of the machine learning problem might be used for investigating the process of spreading a malignant tumor (metastasis) in different organs of a body, provided that real data will be available and appropriate mathematical models will be obtained.
@@
The direction of semisupervised learning we follow here is manifold regularisation, cf. [7], [13], and [50], which attempts to learn the geometry of the input space by exploiting the given unlabeled data. The latter two papers are recent generalisations of the original scalar version of manifold regularisation of [7] to the vector valued setting. In [13], a vector valued version of the graph Laplacian $L$ is used while in [50] $L$ is a general symmetric, positive operator, including the graph Laplacian. The vector valued setting allows one to capture possible dependencies between output variables by the use of, for example, an output graph Laplacian. For a comprehensive discussion on semisupervised learning and a thorough comparison with supervised and unsupervised learning, see the collection [19].
@@
Because reproducing kernel Hilbert spaces make an essential ingredient in these representer type theorems, some historical considerations are in order. The classical article of [2] provides an abstract formalisation of scalar reproducing kernel Hilbert spaces of many previous investigations and applications related to spaces of analytic functions, partial differential equations, and harmonic analysis, as of [47], [8], [52], [11], and [27]. From the point of view of probability theory, [36] investigated linearisations, or feature spaces, associated to scalar positive semidefinite kernels. An equivalent formulation of reproducing kernel Hilbert spaces by Hilbert spaces continuously embeded in a quasicomplete locally convex space was investigated by [63] while a more group theoretical approach was performed in [37] and [38]. Then, scalar valued reproducing kernel Hilbert spaces found many applications in machine learning, see [66] and [61] for a comprehensive list of literature in this direction.
@@
Motivated by problems in operator theory and operator algebras, operator valued positive semidefinite kernels and either their linearisations (Kolmogorov decompositions) or their reproducing kernel Hilbert spaces have been considered by [68], [56], [54], [24], and [20], to cite a few. Investigations of operator valued Hermitian kernels that yield Krein spaces have been performed by [21], [22]. More general operator valued positive semidefinite kernels that yield reproducing kernel VH-spaces have been considered by [26], [3], [4], and [5]. To make things more precise, recall that a vector space ${\mathcal{S}}$ endowed with a map ${\mathcal{S}}\ni s\mapsto s^{*}\in{\mathcal{S}}$ which is conjugate linear and involutive, in the sense that $(s^{*})^{*}=s$ for all $s\in{\mathcal{S}}$, is called a *-vector space. If, in addition, a convex cone ${\mathcal{S}}^{+}$ on ${\mathcal{S}}$ is specified such that for any $s\in{\mathcal{S}}^{+}$ we have $s=s^{*}$, we call ${\mathcal{S}}$ an ordered *-space and, when a complete locally convex topology on ${\mathcal{S}}$ is specified and which is related in a certain fashion with the convex cone ${\mathcal{S}}^{+}$, one calls ${\mathcal{S}}$ an admissible space. The concept of admissible space is a generalisation of the concept of $C^{*}$-algebra which is a mainstream domain in functional analysis, partly due to its strong connections with quantum theories, e.g. see [23] and [10]. Admissible spaces ${\mathcal{S}}$ are then used to consider gramians $[\cdot,\cdot]\colon\mathcal{X}\times{\mathcal{S}}\rightarrow{\mathcal{S}}$, for some vector space $\mathcal{X}$, which are vector valued inner products and that induce a certain topology on $\mathcal{X}$. If this topology is complete then we call $\mathcal{X}$ a VH-space (Vector Hilbert space). The concept of VH-space is a generalisation of the concept of Hilbert modules over $C^{*}$-algebras, e.g. see [41], but it appeared independently and related to problems in probability theory, see [42], [43], and [44]. In case we have a positive semidefinite kernel $K\colon X\times X\rightarrow{\mathcal{S}}$, for some admissible space ${\mathcal{S}}$, one can define a reproducing kernel VH-space as a generalisation of the classical reproducing kernel Hilbert space. For details and many examples see, for example, [4].
@@
We briefly describe the contents of this article. In order to introduce the localised version of the semisupervised, regularised, manifold regularised, and multiview learning problem, in Subsection 2.1 we firstly consider operator valued positive semidefinite kernels for which the entries are localised by a bundle of Hilbert spaces. For these kernels, we show how their reproducing kernel Hilbert spaces are constructed, their relation to the linearisations (Kolmogorov decompositions, feature spaces) and basic properties. Although we have been inspired by the approach in [21], we provide in the appendices proofs for all the statements we make since in this form they cannot be found elsewhere. In [56], reproducing kernel Hilbert spaces are considered for kernels over bundles of locally convex spaces and this might be a very interesting research problem in case applications to machine learning theory might be found.
@@
Further on, in Subsection 3.2 we consider real underlying Hilbert spaces and loss functions that are Gâteaux differentiable and then, in Theorem 3.7 we show that, under this additional assumption, a general representer theorem can be obtained for infinite dimensional input spaces. In Subsection 3.3 we work out the details for the loss functions defined by the least squares which leads to a linear problem in terms of the unkown coefficients, as in [51], while in Subsection 3.4 we work out the details for the loss functions defined by the exponential least squares which lead to a mixed problem, linear and nonlinear. Finally, in Subsection 3.5 we tackle algorithms to obtain approximations of the solutions for the latter machine learning problem by damped Newton approximation methods, as presented in [53], [15], [16], [70]. Since in the framework of systems of nonlinear equations solutions are generally hot topics in current research, much work remains to be done. Some numerical experiments are performed on a toy model for the latter case and the algorithm is tested. We show the robustness of the method on this toy model.
Acknowledgements: The authors thank Naci Saldi and Sorin Costiner for very helpful discussions on available algorithms to solve nonlinear systems of equations and optimisation problems.
@@ 2. A General Semisupervised Regularised and Multiview Machine Learning Problem
Let $X$ be a nonempty set and $\boldsymbol{H}=\{{\mathcal{H}}_{x}\}_{x\in X}$ a bundle of Hilbert spaces over the field ${\mathbb{F}}$, that is either ${\mathbb{R}}$ or ${\mathbb{C}}$, with base $X$, that is, ${\mathcal{H}}_{x}$ is a Hilbert space over ${\mathbb{F}}$, for any $x\in X$. In order to avoid confusion, let us note that the Hilbert spaces ${\mathcal{H}}_{x}$ are actually tagged Hilbert spaces, meaning that for $x\neq y$ the spaces ${\mathcal{H}}_{x}$ and ${\mathcal{H}}_{y}$ are disjoint. An $\boldsymbol{H}$-operator valued kernel $K$ is a mapping defined on $X\times X$ such that $K(y,x)\in{\mathcal{B}}({\mathcal{H}}_{x},{\mathcal{H}}_{y})$ for all $x,y\in X$. Here and throughout this article, if ${\mathcal{H}}$ and ${\mathcal{K}}$ are two Hilbert spaces over the same filed ${\mathbb{F}}$, we denote by ${\mathcal{B}}({\mathcal{H}},{\mathcal{K}})$ the Banach space of all linear and bounded operators $T\colon{\mathcal{H}}\rightarrow{\mathcal{K}}$, endowed with the operator norm. We denote by $\mathfrak{K}(X;\boldsymbol{H})$ the collection of all $\boldsymbol{H}$-operator valued kernels on $X$ and it is clear that $\mathfrak{K}(X;\boldsymbol{H})$ is a vector space over ${\mathbb{F}}$.
@@
In particular, since $\boldsymbol{H}$ consists of tagged Hilbert spaces this means that if either $h\in{\mathcal{H}}_{x}$ or $l\in{\mathcal{H}}_{y}$ are not null and $x\neq y$ then $\widehat{h}\neq\widehat{l}$. However, if $h=0_{{\mathcal{H}}_{x}}$ and $l=0_{{\mathcal{H}}_{y}}$ with $x\neq y$, then $\widehat{h}=\widehat{l}=0_{{\mathcal{F}}(X;\boldsymbol{H})}$. Also, for any $f\in{\mathcal{F}}(X;\boldsymbol{H})$ we have
@@
The following theorem is a general version of some classical results, e.g. [36], [54], [24]. This is also a special case of [21]. We include its proof in the Appendix A.
@@
Also, it is worth mentioning that by (rk2), for each $x\in X$, we actually have a bounded linear operator $K_{x}\colon{\mathcal{H}}_{x}\rightarrow{\mathcal{R}}$ defined by $K_{x}h:=K(\cdot,x)h$, for all $h\in{\mathcal{H}}_{x}$. This operator is bounded, as proven in (4.1). The following result is a generalisation of Moore-Aronszajn Theorem, [52], [2], [48], [18], [17], [50], [35]. Also, it is a special case of [21] and it is proven in the Appendix B. In Appendix C we present a more direct construction of the reproducing kernel Hilbert space induced by an operator valued positive semidefinite kernel.
@@
One of the most important property of a reproducing kernel Hilbert space consists in the fact that, as a function space, its topology makes continuous all evaluation operators, see the proof in the Appendix D.
@@
Secondly, the reproducing kernel Hilbert space ${\mathcal{H}}_{K}$ is associated to a vector valued positive semidefinite kernel for several reasons, but mainly because this is related to the multiview learning, cf. [25], [48], [64], [59], [50], [45], [35], [51], [29], [30], [31], [32]. In this article we consider localised versions of these operator valued reproducing kernel Hilbert spaces that offers flexibility for a larger class of learning problems, as explained in the Introduction, and does not bring additional obstructions, as proven in Subsection 2.1.
@@
$V(y,z)=\frac{1}{1+\exp(z-y)},\quad y,z\in{\mathbb{R}}.$
@@
The third term in (2.15) combines vector valued manifold regularisation, cf. [50], with multiview regularisation, cf. [59] and [67]. The parameter $\gamma_{I}$ may be taken $1$, without loss of generality, since it can be absorbed in $M$. Following [51], the operator multiview regularisation term $\gamma_{I}\langle\boldsymbol{f},M\boldsymbol{f}\rangle_{\boldsymbol{W}^{l+u}}$ is decomposed as
@@
This example follows essentially [51] for a within-view manifold regularisation via multiview graph Laplacians in support vector machine learning, cf. [67]. For manifold regularisation, a data adjaceny graph is defined in such a way that the entries measure the similarity or closeness of pairs of inputs. Given an undirected graph $G=({\mathcal{V}},E)$, where the vertices are ${\mathcal{V}}=\{1,\ldots,n\}$ and edges are simply pairs $(j,k)$, assume that for each edge $(j,k)\in E$ there is a weight $w_{j,k}$, and to each edge $(j,k)\not\in E$ we let $w_{j,k}=0$, in such a way that the weight matrix $W=[w_{j,k}]$ is Hermitian and nonnegative (positive semidefinite).
For example, when each vertex $j$ is associated to a vector $h_{j}\in{\mathbb{R}}^{d}$, we can use the Gaussian weights
@@
In order to simplify the complexity of calculations, cf. [67], for most of the edges $(j,k)$ we take $w_{j,k}=0$ and only for neighbouring $(j,k)$, that is, $\|h_{j}-h_{k}\|_{2}\leq\epsilon$, for some $\epsilon$, we define the weights by (2.17).
Further on, letting $v_{j,j}=\sum_{k=1}^{n}w_{j,k}$ and $v_{j,k}=0$ if $j\neq k$, we make the diagonal matrix $V=[v_{j,k}]$. We work under the assumptions that $v_{j,j}>0$ for all $j=1,\ldots,n$. Then the graph Laplacian matrix is $L:=V-W$, which is positive semidefinite. Sometimes it is useful to work with the normalised graph Laplacian $\widetilde{L}:=V^{-1/2}LV^{-1/2}$.
@@
$\langle\mathbf{f},M_{W}\mathbf{f}\rangle_{{\mathcal{W}}^{u+l}}=\sum_{i=1}^{m}\sum_{j,k=1,\ j<k}^{l+u}w_{j,k}^{i}\|f^{i}(x_{j})-f^{i}(x_{k})\|^{2}_{\mathcal{Y}},$
@@
$E_{\boldsymbol{C},\boldsymbol{x}}f=\left(C_{x_{1}}K_{x_{1}}^{*}f,\dots,C_{x_{l}}K_{x_{l}}^{*}f\right)=\left(C_{x_{1}}f(x_{1}),\dots,C_{x_{l}}f(x_{l})\right),\quad f\in{\mathcal{H}}_{K}.$ (2.21)